# Consistency of EFT illuminated via relative entropy: A case study in scalar field theory

### v1

@@ Abstract
Relative entropy is a non-negative quantity and offers a powerful means of achieving a unified understanding of fundamental properties in physics, including the second law of thermodynamics and positivity bounds on effective field theories (EFTs). We analyze the relative entropy in scalar field theories and show that the non-negativity of relative entropy is potentially violated in perturbative calculations based on operator and loop expansions. Conversely, this suggests that the consistency of the EFT description in the scalar field theory can be identified by the sign of the relative entropy. In fact, we revisit an EFT of single-field inflation and present a relation between its non-linear parameter $f_{\rm NL}$ and the consistency condition of the EFT description derived from the relative entropy method. We find that interesting regions of $f_{\rm NL}$ that are observationally allowed can be constrained from the relative entropy by imposing the consistency of the EFT description.
@@ 1 Introduction
The instances of success mentioned above are grounded in the unitarity of the $S$-matrix. However, the unitarity of physical theories could manifest in quantities other than the $S$-matrix. A new method Cao:2022iqh ; Cao:2022ajt has been suggested to evaluate the consistency of physical quantities, such as Wilson coefficients of EFTs, with the existence of the UV theory. This method focuses on the non-negative relative entropy 10.1214/aoms/1177729694 ; 10.2996/kmj/1138844604 ; RevModPhys.50.221 that arises from the unitarity of the theory. The relative entropy, also known as Kullback-Leibler divergence, is a non-negative quantity defined by two probability distribution functions. It is zero only when the two probability distribution functions are equal. The relative entropy, in brief, measures the disparity between two probability distributions. In Refs. Cao:2022ajt ; Cao:2022iqh , it was suggested that the relative entropy can be utilized to quantify the disparity between theories with and without interactions between heavy and light degrees of freedom, i.e., a transfer of UV information to EFTs consisting of light degrees of freedom through the interactions. There, a canonical distribution function is considered for the probability distribution function of the theory. In this case, the non-negativity of relative entropy holds when the Hamiltonian of systems respects Hermiticity, i.e., unitarity for the time evolution of systems. Refs. Cao:2022ajt ; Cao:2022iqh indicate that the non-negativity of relative entropy and the consideration of the $S$-matrix lead to several common outcomes, such as positivity bounds on the SMEFT gauge bosonic operator Remmen:2019cyz and adherence to the mild weak gravity conjecture Vafa:2005ui ; Arkani-Hamed:2006emk ; Kats:2006xp ; Cheung:2018cwt ; Cheung:2019cwi ; Loges:2019jzs ; Reall:2019sah ; Goon:2019faz ; Bellazzini:2019xts ; Hamada:2018dde ; Arkani-Hamed:2021ajd . One notable aspect of this relative entropy consideration is that it presents constraints on the EFT in a different approach than the $S$ matrix consideration. This motivates applying entropy considerations to various phenomena to reveal possible connections and differences between relative entropy considerations and $S$-matrix considerations and to understand when relative entropy methods can have more vital implications.
@@
For several reasons, EFTs consisting of scalar fields could be candidates for adapting the relative entropy method. The first reason is that while scalar field theories are relevant to exciting phenomena such as inflation and phase transitions, considerations of relative entropy have not yet been comprehensively adapted. As will be explained later, the EFT consisting of scalar fields provides an excellent example of considering a relation between the effectiveness of perturbation calculations and the unitarity appearing as non-negativity of relative entropy, as the tree-level effects may be more dominant in the corrections from UV theory to the EFT. In the scattering amplitude considerations, the breakdown of perturbation calculations (where the validity of such calculations is often judged by the negligible contribution of loop-level terms relative to tree-level terms) can lead to a violation of inequalities stemming from the unitarity. However, the relation between the breakdown of the non-negativity of relative entropy and the breakdown of perturbation calculations is still unclear. It will be essential to clarify the relation and its phenomenological consequences in the theory of scalar fields. Second, inflation models described by EFTs involving scalar fields would provide a valuable example to validate the applicability of relative entropy calculated for the background field, especially since the inflationary background field violates Lorentz invariance. For instance, in conventional positivity bounds, Lorentz symmetry is an essential factor in determining the behavior of the scattering amplitude, and several techniques have been made in previous studies to derive bounds on the inflationary EFT Baumann:2015nta ; Grall:2021xxm ; Freytsis:2022aho . On the other hand, the Lorentz invariance is not an essential factor in the relative entropy method, as it has been adapted to non-relativistic spin systems, for instance. Hence, the relative entropy method might be more potent in inflation models. Third, it has been pointed out that in a single-field inflationary EFT, assuming a particular class of UV theories Kim:2021pbr (two scalar field theories), the scattering amplitudes at high energy scales violate the unitarity. The single-field inflationary EFT model Garriga:1999vw ; Chen:2006nt is one of the attractive models because of its potential to generate observable non-Gaussianity. On the other hand, in Ref. Kim:2021pbr , it is pointed out that if one considers the two scalar field theories as UV theories of their EFTs, the non-Gaussianity of $|f_{\rm NL}|\gtrsim 1$ can only be realized by turning on interactions that violate the unitarity at high energy scales. Revisiting this point with considerations of relative entropy would lead to understanding the relation and differences between considerations of relative entropy and perturbative unitarity based on scattering amplitudes. For the above reasons, state-of-the-art research would encourage considering entropy in scalar field EFTs, including inflation models.
@@
This paper examines the role of relative entropy in scalar field theories with and without interactions between heavy and light fields. We will demonstrate that the non-negativity of relative entropy can be violated in perturbative calculations based on operator and loop expansions. As a byproduct, we highlight that the relative entropy can be utilized to identify theories for which the perturbation calculations are invalid. As will be explained, the EFT expansions are valid for EFTs derived from weakly coupled theories, and the sign of the relative entropy can help determine the theories for which the EFT description is valid. When we apply this concept to the two-scalar inflation in two typical cases: the U(1) symmetry potential and the U(1) symmetry-breaking potential, we observe that the U(1) symmetry case represents a weakly coupled theory, while the U(1) symmetry-breaking case can be seen as a strong coupling theory. Furthermore, we provide constraints on the Wilson coefficients, focusing on higher-dimensional operators in a specific EFT of a single scalar field derived from general weakly coupled theories. Comparing these constraints with the parameter domain suggested by the non-linear parameter $f_{\rm NL}$ allowed by observations Planck:2019kim , we argue that the sign of the non-linear parameter $f_{\rm NL}$ can be severely constrained under the valid EFT description.
@@
The rest of the paper is organized as follows. In Sec. 2, we briefly review Refs. Cao:2022iqh ; Cao:2022ajt and then illustrate how to calculate the relative entropy in three typical cases of scalar field theories. In particular, the definition of the weakly coupled theory related to the valid EFT description is presented in Sec. 2.2.1. In Sec. 3, we apply the results of Sec. 2 to an EFT of single-field inflation, and consequences of the non-negativity of relative entropy are examined numerically and analytically in three UV theories, including a general theory, with an emphasis on the validity of the EFT description. In Sec. 4, as a phenomenological application example of results of Sec. 2, we present a relation between the non-linear parameter in the EFT of single-field inflation and the parameter regions where the relative entropy indicates the validity of the EFT description. In Sec. 5, assumptions related to the definition of the weakly coupled theory are discussed. The main findings and outlook of our analysis are summarized in Sec. 6. The Appendix contains procedures of field redefinitions to make calculations easier (Appendix A) and the calculations of the effective actions up to the sixth order of the perturbation expansion (Appendix B), as well as a relation between different bases of field (Appendix C).
@@ 2 Preliminaries
We will commence with a brief review of Refs. Cao:2022iqh ; Cao:2022ajt to enhance our understanding of the forthcoming sections. Then, we will illustrate how to calculate the relative entropy in theories where perturbative calculations are valid (detailed conditions of the valid perturbations are summarized in Sec. 2.2.1) with a specific focus on scalar field theories. Drawing upon the results and employing a methodology based on relative entropy Cao:2022iqh ; Cao:2022ajt , we aim to assess whether the EFT originates from a UV theory and to determine the parameter regions within which the EFT description remains valid.
@@ 2.1 Non-negativity of relative entropy arising in unitary theories
This parameter $\lambda$ is introduced to explicitly specify the order of perturbation expansion in the interaction $I_{\rm I}$, and therefore, it is permissible to select $\lambda=1$ subsequent to the calculation of the relative entropy. The information from the UV theory is conveyed to an EFT through the interaction $I_{\rm I}$ by integrating $\Phi$, and the disparity between theories with and without the interaction $I_{\rm I}$ signifies the transferred information. In Refs Cao:2022ajt ; Cao:2022iqh , the relative entropy between two theories is calculated for two probability distribution functions, which are defined as follows:
@@
where the path integral for $\phi$ is performed if $\phi$ is dynamical, but not in the case of an external field (please look at Sec. 2.2 for specific details). In this context, if the theory complies with unitarity, i.e., if the actions $I_{0}$ and $I_{\text{I}}$ are Hermitian, and if considerations of a low energy scale, such as heavy degrees of freedom not being dynamical, are taken into account, the conditions $P_{\rm R}=P_{\rm R}^{\dagger}$ and $P_{\rm T}=P_{\rm T}^{\dagger}$ are satisfied. The non-negativity of relative entropy then follows from this premise. In short, the non-negativity of relative entropy under the probability distribution functions (5) corresponds to the concept of unitarity. Considering the significance of this point, we will examine this behavior from a quantum mechanical standpoint later.
@@
where $W_{0}:=-\ln Z_{0}$ and $W_{\lambda}:=-\ln Z_{\lambda}$ represent Euclidean effective actions, and we performed several straightforward calculations, $\ln P_{\rm R}=-I_{0}-\ln Z_{0}$, $\ln P_{\rm T}=-I_{\lambda}-\ln Z_{\lambda}$, $\lim_{\lambda\to 0}P_{\rm R}=P_{\rm T}$, and used the non-negativity of relative entropy (2). The last term in eq. (7) is a partial derivative concerning $\lambda$ while keeping $\phi$ and $\Phi$ constant (several examples will be presented in Sec. 2.2). From eq. (7), it was found that the relative entropy can be expressed in terms of the effective actions, and the calculation of the relative entropy, therefore, reduces to the calculation of the effective action. In the relative entropy calculations of this work, we assume low-energy regions where heavy degrees of freedom are non-dynamical. Consequently, the effective action $W_{\lambda}$ indicates the EFT corresponding to $I_{\lambda}$ within the low-energy domain. In light of this observation, the inequality (7) means that the low-energy EFT of the theory $I_{\lambda}$ must maintain a specific relation (please look at Refs. Cao:2022ajt ; Cao:2022iqh and Sec. 3 for further details). In the regime of perturbative theories, each term of the EFT can be determined by, e.g., matching the scattering amplitudes, and similarly, the relative entropy is also calculable perturbatively. The effective actions within this work are calculated under specific external background fields or light field solutions, as will be demonstrated later with several examples. In Sec. 3, we will calculate the effective actions within an inflationary background and examine the implications of eq. (7) in a specific class of EFT of inflation.
@@
In the derivation above, we utilized $\rho_{\rm R}^{\dagger}=\rho_{\rm R}$, $\rho_{\rm T}^{\dagger}=\rho_{\rm T}$, and ${\rm Tr}\,[\rho_{\rm R}]={\rm Tr}\,[\rho_{\rm T}]=1$. Although the derivation of this inequality (9) is a reformulation of eq. (7), it would be more explicit that the non-negativity of relative entropy represents the unitarity.
It is also worth mentioning that the conditions of the probability distribution function ${\rm Tr}\,[\rho_{\rm R}]={\rm Tr}\,[\rho_{\rm T}]=1$ must be maintained consistently over time due to the conservation of probability through the unitarity. Let us consider the quantum state of the interacting theory at time zero, denoted as $\rho_{\rm T}(0):=\rho_{\rm T}$. In this case, the quantum state of this interacting theory at time $t$ becomes $\rho_{\rm T}(t)=e^{-iH_{\lambda}t}\rho_{\rm T}(0)e^{+iH_{\lambda}^{\dagger}t}$. The time evolution of this system described by $H_{\rm T}$ is elaborated as follows:
@@
where we defined $h_{\rm T}:=(H_{\rm T}+H_{\rm T}^{\dagger})/2$, and $\gamma_{\rm T}:=i(H_{\rm T}-H_{\rm T}^{\dagger})/2$. In the context of theories respecting the unitarity, $h_{\rm T}=H_{\rm T}$ and $\gamma_{\rm T}=0$ hold, resulting in the right-hand side of eq (10) becoming zero. The same discussions apply even to the system defined by $H_{\rm R}$. In unitary theories, the condition ${\rm Tr}\,[\rho_{\rm R}]={\rm Tr}\,[\rho_{\rm T}]=1$ is thus upheld consistently over time. These arguments further demonstrate that the non-negativity of relative entropy reflects the unitary time evolution. As studied in Refs. Cao:2022ajt ; Cao:2022iqh , the non-negativity of relative entropy, which signifies the unitarity, is linked to fundamental physics properties encompassing the second law of thermodynamics, positivity bounds on EFTs, and the mild weak gravity conjecture. Any violation of this non-negativity results in the emergence of abnormal behavior within the theory. In the forthcoming sections, we will explore an aspect of the non-negative relative entropy that assesses the validity of perturbations utilized in calculating effective actions.

### v5

@@ Abstract
Relative entropy is a non-negative quantity and offers a powerful means of achieving a unified understanding of fundamental properties in physics, including the second law of thermodynamics and positivity bounds on effective field theories (EFTs). We analyze the relative entropy in scalar field theories and show that the non-negativity of relative entropy is potentially violated in perturbative calculations based on operator and loop expansions. Conversely, this suggests that the consistency of the EFT description in the scalar field theory can be identified by the sign of the relative entropy. In fact, we revisit an EFT of single-field inflation and present a relation between its non-linear parameter $f_{\rm NL}$ and the consistency condition of the EFT description derived from the relative entropy method. We find that interesting regions of $f_{\rm NL}$ that are observationally allowed can be constrained from the relative entropy by imposing the consistency of the EFT description when the EFT is generated via the interaction with heavy fields in UV theories.
@@ 1 Introduction
The instances of success mentioned above are grounded in the unitarity of the $S$-matrix. However, the unitarity of physical theories could manifest in quantities other than the $S$-matrix. A new method Cao:2022iqh ; Cao:2022ajt has been proposed to assess the consistency of physical quantities, such as Wilson coefficients of EFTs, with underlying UV theories. This method focuses on the non-negative relative entropy 10.1214/aoms/1177729694 ; 10.2996/kmj/1138844604 ; RevModPhys.50.221 that arises from the unitarity of the theory. The relative entropy, also known as Kullback-Leibler divergence, is a non-negative quantity defined by two probability distribution functions. It is zero only when the two probability distribution functions are equal. The relative entropy, in brief, measures the disparity between two probability distributions. In Refs. Cao:2022ajt ; Cao:2022iqh , it was suggested that the relative entropy can be utilized to quantify the disparity between theories with and without interactions between heavy and light degrees of freedom, i.e., a transfer of UV information to EFTs consisting of light degrees of freedom through the interactions. There, a canonical distribution function is considered for the probability distribution function of the theory. In this case, the non-negativity of relative entropy holds when the Hamiltonian of systems respects Hermiticity, i.e., unitarity for the time evolution of systems. Refs. Cao:2022ajt ; Cao:2022iqh indicate that the non-negativity of relative entropy and the consideration of the $S$-matrix lead to several common outcomes, such as positivity bounds on the SMEFT gauge bosonic operator Remmen:2019cyz and adherence to the mild weak gravity conjecture Vafa:2005ui ; Arkani-Hamed:2006emk ; Kats:2006xp ; Cheung:2018cwt ; Cheung:2019cwi ; Loges:2019jzs ; Reall:2019sah ; Goon:2019faz ; Bellazzini:2019xts ; Hamada:2018dde ; Arkani-Hamed:2021ajd ; Bittar:2024xuc . One notable aspect of this relative entropy consideration is that it presents constraints on the EFT in a different approach than the $S$ matrix consideration. This motivates applying entropy considerations to various phenomena to reveal possible connections and differences between relative entropy considerations and $S$-matrix considerations and to understand when relative entropy methods can have more vital implications.
@@
For several reasons, EFTs consisting of scalar fields could be candidates for adapting the relative entropy method. The first reason is that while scalar field theories are relevant to exciting phenomena such as inflation and phase transitions, considerations of relative entropy have not yet been comprehensively adapted. As will be explained later, the EFT consisting of scalar fields provides an excellent example of considering a relation between the effectiveness of perturbative calculations and the unitarity appearing as non-negativity of relative entropy, because the tree-level effects may be more dominant in the corrections from UV theory to the EFT [^1] For instance, loop-level effects may dominate in the EFT derived from a UV theory, where fermion fields are the only heavy degrees of freedom. However, if heavy scalar fields are present in the UV theory, tree-level effects from heavy scalar fields dominate under naive loop expansions. See Sec. 2.2 for details. . In the scattering amplitude considerations, the breakdown of perturbative calculations (where the validity of such calculations is often judged by the negligible contribution of loop-level terms relative to tree-level terms) can lead to a violation of inequalities stemming from the unitarity. However, the relation between the breakdown of the non-negativity of relative entropy and the breakdown of perturbative calculations is still unclear. It will be essential to clarify the relation and its phenomenological consequences in the theory of scalar fields. Second, inflation models described by EFTs involving scalar fields would provide a valuable example to validate the applicability of relative entropy calculated for the background field, especially since the inflationary background field violates Lorentz invariance. For instance, in conventional positivity bounds, Lorentz symmetry is an essential factor in determining the behavior of the scattering amplitude, and several techniques have been made in previous studies to derive bounds on the inflationary EFT Baumann:2015nta ; Grall:2021xxm ; Freytsis:2022aho . On the other hand, the Lorentz invariance is not an essential factor in the relative entropy method, as it has been adapted to non-relativistic spin systems, for instance. Hence, the relative entropy method might be more potent in inflation models [^2] The purity for a cosmological model has also been studied in past works, e.g., Ref. Colas:2024ysu .. Third, it has been pointed out that in a single-field inflationary EFT, assuming a particular class of UV theories Kim:2021pbr (two scalar field theories), the scattering amplitudes at high energy scales violate the unitarity. The single-field inflationary EFT model Garriga:1999vw ; Chen:2006nt is one of the attractive models because of its potential to generate observable non-Gaussianity. On the other hand, it is pointed out that if one considers the two scalar field theories as UV theories of their EFTs, the non-Gaussianity of $|f_{\rm NL}|\gtrsim 1$ can only be realized by turning on interactions that violate the unitarity at high energy scales. Revisiting these point with considerations of relative entropy would lead to understanding the relation and differences between considerations of relative entropy and perturbative unitarity based on scattering amplitudes. For the above reasons, state-of-the-art research would encourage considering entropy in scalar field EFTs, including inflation models.
@@
This paper examines the role of relative entropy in scalar field theories with and without interactions between heavy and light fields. We will demonstrate that the non-negativity of relative entropy can be violated in perturbative calculations based on operator and loop expansions. As a byproduct, we highlight that the relative entropy can be utilized to identify theories for which the perturbative calculations are invalid. As will be explained, the EFT expansions are valid for EFTs derived from weakly coupled theories, and the sign of the relative entropy can help determine the theories for which the EFT description is valid. When applying this concept to two-scalar inflation, we consider two typical cases: one with a renormalizable potential respecting U(1) symmetry and another with a modified potential. We observe that the former U(1)-symmetric case corresponds to a weakly coupled theory, whereas the modified case can be interpreted as a strongly coupled one. Furthermore, we provide constraints on the Wilson coefficients, focusing on higher-dimensional operators in a specific EFT of a single scalar field derived from general weakly coupled theories. Comparing these constraints with the parameter domain suggested by the non-linear parameter $f_{\rm NL}$ allowed by observations Planck:2019kim , we argue that the sign of the non-linear parameter $f_{\rm NL}$ can be severely constrained under the valid EFT description.
@@
The rest of the paper is organized as follows. In Sec. 2, we briefly review Refs. Cao:2022iqh ; Cao:2022ajt and then illustrate how to calculate the relative entropy in three typical cases of scalar field theories. In particular, the definition of the weakly coupled theory related to the valid EFT description is presented in Sec. 2.2. In Sec. 3, we apply the results of Sec. 2 to an EFT of single-field inflation, and consequences of the non-negativity of relative entropy are examined numerically and analytically in three UV theories, including a general theory, with an emphasis on the validity of the EFT description. In Sec. 4, as a phenomenological application example of results of Sec. 2, we present a relation between the non-linear parameter in the EFT of single-field inflation and the parameter regions where the relative entropy indicates the validity of the EFT description. The main findings and outlook of our analysis are summarized in Sec. 5. The Appendix contains procedures of field redefinitions to make calculations easier (Appendix A), perturbative calculations of $\widetilde{\phi}_{1,\lambda}$ (Appendix B), the calculations of the effective actions up to the sixth order of the perturbative expansion (Appendix C), and a derivation of eq. (48) (Appendix D), as well as a relation between different bases of field (Appendix E).
@@ 2 Preliminaries
We will commence with a brief review of Refs. Cao:2022iqh ; Cao:2022ajt to enhance our understanding of the forthcoming sections. Then, we will illustrate how to calculate the relative entropy in theories where perturbative calculations are valid (detailed conditions of the valid perturbation theory are summarized in Sec. 2.2) with a specific focus on scalar field theories. Drawing upon the results and employing a methodology based on relative entropy Cao:2022iqh ; Cao:2022ajt , we aim to assess whether the EFT originates from a UV theory and to determine the parameter regions within which the EFT description remains valid.
@@ 2.1 Non-negativity of relative entropy arising in unitary theories
This parameter $\lambda$ is introduced to explicitly specify the order of perturbative expansion in the interaction $I_{\rm I}$, and therefore, it is permissible to select $\lambda=1$ subsequent to the calculation of the relative entropy. The information from the UV theory is conveyed to an EFT through the interaction $I_{\rm I}$ by integrating $\Phi$, and the disparity between theories with and without the interaction $I_{\rm I}$ signifies the transferred information. In Refs Cao:2022ajt ; Cao:2022iqh , the relative entropy between two theories is calculated for two probability distribution functions, which are defined as follows:
@@
where the path integral for $\phi$ is performed if $\phi$ is dynamical, but not in the case of an external field (please look at Sec. 2.3 for specific details). In this context, if the theory complies with unitarity, i.e., if the actions $I_{0}$ and $I_{\text{I}}$ are Hermitian, and if considerations of a low energy scale, such as heavy degrees of freedom not being dynamical [^3] After integrating out those heavy fields, the effects of dynamical heavy fields can manifest as an imaginary contribution to the effective action at high energy scales, potentially violating the non-negativity condition of relative entropy. To avoid these effects, we focus on a low energy regime where those heavy fields are no longer dynamical. , are taken into account, the conditions $P_{\rm R}=P_{\rm R}^{\dagger}$ and $P_{\rm T}=P_{\rm T}^{\dagger}$ are satisfied. The non-negativity of relative entropy then follows from this premise. In short, the non-negativity of relative entropy under the probability distribution functions (5) corresponds to the concept of unitarity. Considering the significance of this point, we will examine this behavior from a quantum mechanical standpoint later.
@@
where $W_{0}:=-\ln Z_{0}$ and $W_{\lambda}:=-\ln Z_{\lambda}$ represent Euclidean effective actions, and we performed several straightforward calculations, $\ln P_{\rm R}=-I_{0}-\ln Z_{0}$, $\ln P_{\rm T}=-I_{\lambda}-\ln Z_{\lambda}$, $\lim_{\lambda\to 0}P_{\rm R}=P_{\rm T}$, and used the non-negativity of relative entropy (2). The last term in eq. (7) is a partial derivative concerning $\lambda$ while keeping $\phi$ and $\Phi$ constant (several examples will be presented in Sec. 2.3 and Appendix D). From eq. (7), it was found that the relative entropy can be expressed in terms of the effective actions, and the calculation of the relative entropy, therefore, reduces to the calculation of the effective action. In the relative entropy calculations of this work, we assume low-energy regions where heavy degrees of freedom are non-dynamical. Consequently, the effective action $W_{\lambda}$ indicates the EFT corresponding to $I_{\lambda}$ within the low-energy domain. In light of this observation, the inequality (7) means that the low-energy EFT of the theory $I_{\lambda}$ must maintain a specific relation (please look at Refs. Cao:2022ajt ; Cao:2022iqh and Sec. 3 for further details). In the regime of perturbative theories, each term of the EFT can be determined by, e.g., matching the scattering amplitudes, and similarly, the relative entropy is also calculable perturbatively. The effective actions within this work are calculated under specific external background fields or light field solutions, as will be demonstrated later with several examples. In Sec. 3, we will calculate the effective actions within an inflationary background and examine the implications of eq. (7) in a specific class of EFT of inflation.
@@
In the derivation above, we utilized $\rho_{\rm R}^{\dagger}=\rho_{\rm R}$, $\rho_{\rm T}^{\dagger}=\rho_{\rm T}$, and ${\rm Tr}\,[\rho_{\rm R}]={\rm Tr}\,[\rho_{\rm T}]=1$. Although the derivation of this inequality (9) is a reformulation of eq. (7), it becomes more explicit that the non-negativity of relative entropy represents the unitarity.
It is also worth mentioning that the conditions of the probability distribution function ${\rm Tr}\,[\rho_{\rm R}]={\rm Tr}\,[\rho_{\rm T}]=1$ must be maintained consistently over time due to the conservation of probability through the unitarity. Let us consider the quantum state of the interacting theory at time zero, denoted as $\rho_{\rm T}(0):=\rho_{\rm T}$. In this case, the quantum state of this interacting theory at time $t$ becomes $\rho_{\rm T}(t)=e^{-iH_{\lambda}t}\rho_{\rm T}(0)e^{+iH_{\lambda}^{\dagger}t}$. The time evolution of this system described by $H_{\lambda}$ is elaborated as follows:
@@
where we defined $h_{\rm T}:=(H_{\lambda}+H_{\lambda}^{\dagger})/2$, and $\gamma_{\rm T}:=i(H_{\lambda}-H_{\lambda}^{\dagger})/2$. In the context of theories respecting the unitarity, $h_{\rm T}=H_{\lambda}$ and $\gamma_{\rm T}=0$ hold, resulting in the right-hand side of eq (10) becoming zero. The same discussions apply even to the system defined by $H_{0}$. In unitary theories, the condition ${\rm Tr}\,[\rho_{\rm R}]={\rm Tr}\,[\rho_{\rm T}]=1$ is thus upheld consistently over time. These arguments further demonstrate that the non-negativity of relative entropy reflects the unitary time evolution. As studied in Refs. Cao:2022ajt ; Cao:2022iqh , the non-negativity of relative entropy, which signifies the unitarity, is linked to fundamental physics properties encompassing the second law of thermodynamics, positivity bounds on EFTs, and the mild weak gravity conjecture. Any violation of this non-negativity results in the emergence of abnormal behavior within the theory. In the forthcoming sections, we will explore an aspect of the non-negative relative entropy that assesses the validity of perturbative calculations utilized in calculating effective actions.
2.2 Weakly coupled theory
Now, we elucidate assumptions in calculating the effective action as fundamental components of the relative entropy. In subsequent sections, in the perturbative calculations of the EFT $W_{\lambda}$, it is assumed that the interaction $I_{\rm I}$ is small and that the following two conditions are upheld: {itembox}[c] Weakly coupled theory
These two assumptions imply that the interaction between heavy and light fields is weak, and consequently, we refer to theories that satisfy these assumptions as weakly coupled theories. In contrast, throughout this work, theories that do not satisfy one of the two assumptions are referred to as strongly coupled theories. The first condition above is often assumed in the context of perturbative unitarity of $S$-matrix, and exploring parameter regions of theory satisfying its validity is of significant importance, e.g., in predicting the SM Higgs mass Lee:1977eg ; PhysRevLett.38.883 . It is well-known that perturbative expansions about the interaction arise during the calculation of scattering amplitudes, and the effective action can be obtained by matching these scattering amplitudes. Based on this perspective, the tree-level approximation utilized in perturbative unitarity within the $S$-matrix is similar to the tree-level approximation applied in the EFT above calculations. On the other hand, the second condition represents the validity of operator expansion, meaning the validity of the EFT description in low-energy regions. If the operator expansion is invalid, the phenomenon cannot be adequately described without adding an infinite number of operators. The determination of the Wilson coefficients of an infinite number of operators requires a UV theory that corresponds to the EFT. We thus regard any theory for which the operator expansion is invalid as unsuitable for the EFT description.
In the subsequent sections, we will not consider the loop-level contributions in evaluating the relative entropy because the emphasis is on its non-negativity in the weakly coupled theory. For instance, if the UV theory includes heavy fermions in addition to scalar fields, assuming the weakly coupled theory means that loop-level contributions arising from the fermion fields are negligible. However, our results of this study might not be applicable when the main contribution in the EFT is at the loop-level, such that the heavy degrees of freedom in the UV theory consist only of fermion fields. As studied in Ref. Cao:2022ajt ; Cao:2022iqh , there are classes of UV theories to which the results of this study can be adapted at the loop-level, while there are also classes of UV theories to which they cannot be adapted [^4] More specifically, there are theories for which the third term in the last line of eq. (7) is not zero at the loop-level. Such theories are not considered in this study, and the results of this study may not be available.. The calculations of the relative entropy at the loop-level have been studied in Ref. Cao:2022ajt ; Cao:2022iqh ; please look at the details there. Assuming the weakly coupled theory also implies the validity of saddle-point approximation in the path integral, as mentioned in Ref. Cao:2022ajt ; Cao:2022iqh . This is because classical-level evaluations are inadequate when performing Euclidean path integrals near maxima points of the action, and higher-order loop-level corrections are essential.
Regarding the second condition of the weakly coupled theory, we set the condition for the validity of the operator expansion to be able to truncate higher-order operators at arbitrary orders quantitatively. For instance, the perturbative calculation in the following case is valid, although, at first glance, it does not meet this validity condition. Suppose that, up to a particular order $n$-th of the operator, the higher-order contributions of the operators are more significant than the lower-order contributions, but when the order is bigger than $n$-th, the contributions of the higher-order operators are sufficiently smaller than those of the lower-order operators. In this case, the higher-order operator cannot be quantitatively truncated at any order smaller than $n$-th, but the perturbative calculation can be regarded as working well for sufficiently high-order perturbative calculations. However, the perturbative calculation in this case would not be valid because it is essential to evaluate them to all orders, as the specific order $n$ can only be determined by considering the contributions from all perturbative orders. In the present work, the weakly coupled theory is defined so that situations requiring infinite-order evaluation of the perturbative expansion are not included in the valid perturbation regions. Given these assumptions in the perturbative calculation, we now turn to the relative entropy in scalar field theories.